# Affordable Generative Agents

### v1

@@ B.2 Distinguishing Criteria
Appendix C The LLM Framework of VirtualHome
VirtualHome is a 3D simulated domestic environment including 115 items with varying properties and states. We’ve developed an LLM agent programmed to simulate a fulfilling day at home. Unlike other existing approaches, we do not test on fixed tasks, but rather aim for LLM agents to generate believable activities and decompose them into detailed actions. The implementation is depicted in Figures 7 to 9, with components triggering the LLM highlighted in blue.
@@ Appendix C The LLM Framework of VirtualHome
Appendix D An Effective Way for Emergence of Diverse Social Behavior
LLMs are pretrained on extensive text data to predict subsequent tokens and refined via reinforcement learning with human feedback (RLHF) to align with human preferences. This ensures that LLMs respond in accordance with the instructions and requirements of the given prompt. For instance, in Generative Agents (Park et al., 2023), the agent, Isabella Rodriguez, is designed to be a coffee shop manager intending to host a Valentine’s Day party. In the simulated town, Isabella repeatedly engages in activities such as managing the coffee shop, serving customers, and informing them about the upcoming Valentine’s event. However, in the real world, human behavior patterns are not so predictable. Human actions and thoughts are influenced not only by the surrounding environment but also by spontaneous disturbances. In the field of cognitive science, this is known as mind wandering, a phenomenon in which where humans experience task-unrelated or stimulus-unrelated thoughts (Seli et al., 2016; Christoff et al., 2016). The strong consistency inherent in LLMs constrains the emergence of the agents’ diverse behavior.
Besides, the agent framework also impacts the agent’s behavior. Generative Agents emphasizes the importance of memory in constructing self-consistent and believable agents, incorporating a long-term memory module and a memory retrieval model. The long-term memory module maintains a comprehensive record of the agent’s experiences. Due to the limited length of prompts, not all events can be included in the prompt as part of the memory. Thus, a retrieval model is designed to take current events as input and return a subset of memory events. The retrieval model judges based on three criteria: 1) Recency, where more recent events score higher, 2) Importance, where events are scored for significance based on the LLM’s interpretation, and 3) Relevance, scored through cosine similarity to determine the similarity of events. We argue that this framework ensures strong consistency in the agent but results in agent’s limited behavior. We analyze Isabella Rodriguez’s memory events sampled from one experiment, filtering out ”idle” events, leaving 740, of which 506 were related to the profile setup. Furthermore, we employed the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) to analyze the frequency of events. Table 4 presents the five most and least frequent events in terms of cluster size, revealing that the most frequent events largely align with the agent’s inner profile setup.
@@ Appendix D An Effective Way for Emergence of Diverse Social Behavior
To ensure that agent’s actions align with its inner profile while avoiding purely mechanical behavior, we incorporate an additional module into the memory design of Generative Agents, which we call ”mind wandering.” In addition to sampling highly relevant events to assist the agent in making accurate decisions, we randomly sample unrelated events to influence the agent’s behavior. For LLMs, we include random content in the prompt, which will have a certain impact on the output response. For the agent, we want it to behave like a real person, where spontaneous thoughts may influence its decision-making, or leading to a shift in the conversation topic. In our implementation, to prevent the frequent sampling of events that match the agent’s inner profile, we conduct weighted sampling based on the repetition rate. We employ the DBSCAN algorithm to cluster the embedding features of agent’s memory events $X=\{x_{1},x_{2},...,x_{n}\}$ into $k$ clusters denoted as $\{C_{1},C_{2},...,C_{k}\}$. For each cluster $C_{i}$, where $i=1,2,...,k$, we compute the number of events, $|C_{i}|$, contained within. The sampling probability $p_{i}$ for events in cluster $C_{i}$ is defined as:

### v2

@@ D.2 Distinguishing Criteria
Appendix E The LLM Framework of VirtualHome
VirtualHome is a 3D simulated domestic environment including 115 items with varying properties and states. We’ve developed an LLM agent programmed to simulate a fulfilling day at home. Unlike other existing approaches, we do not test on fixed tasks, but rather aim for LLM agents to generate believable activities and decompose them into detailed actions. The implementation is depicted in Figures 8 to 10, with components triggering the LLM highlighted in blue.
@@ Appendix E The LLM Framework of VirtualHome
Appendix F An Effective Way for Emergence of Diverse Social Behavior
LLMs are pretrained on extensive text data to predict subsequent tokens and refined via reinforcement learning with human feedback (RLHF) to align with human preferences. This ensures that LLMs respond in accordance with the instructions and requirements of the given prompt. For instance, in Generative Agents Park et al. (2023), the agent, Isabella Rodriguez, is designed to be a coffee shop manager intending to host a Valentine’s Day party. In the simulated town, Isabella repeatedly engages in activities such as managing the coffee shop, serving customers, and informing them about the upcoming Valentine’s event. However, in the real world, human behavior patterns are not so predictable. Human actions and thoughts are influenced not only by the surrounding environment but also by spontaneous disturbances. In the field of cognitive science, this is known as mind wandering, a phenomenon in which where humans experience task-unrelated or stimulus-unrelated thoughts Seli et al. (2016); Christoff et al. (2016). The strong consistency inherent in LLMs constrains the emergence of the agents’ diverse behavior.
Besides, the agent framework also impacts the agent’s behavior. Generative Agents emphasizes the importance of memory in constructing self-consistent and believable agents, incorporating a long-term memory module and a memory retrieval model. The long-term memory module maintains a comprehensive record of the agent’s experiences. Due to the limited length of prompts, not all events can be included in the prompt as part of the memory. Thus, a retrieval model is designed to take current events as input and return a subset of memory events. The retrieval model judges based on three criteria: 1) Recency, where more recent events score higher, 2) Importance, where events are scored for significance based on the LLM’s interpretation, and 3) Relevance, scored through cosine similarity to determine the similarity of events. We argue that this framework ensures strong consistency in the agent but results in agent’s limited behavior. We analyze Isabella Rodriguez’s memory events sampled from one experiment, filtering out "idle" events, leaving 740, of which 506 were related to the profile setup. Furthermore, we employed the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) to analyze the frequency of events. Table 6 presents the five most and least frequent events in terms of cluster size, revealing that the most frequent events largely align with the agent’s inner profile setup.
@@ Appendix F An Effective Way for Emergence of Diverse Social Behavior
To ensure that agent’s actions align with its inner profile while avoiding purely mechanical behavior, we incorporate an additional module into the memory design of Generative Agents, which we call "mind wandering." In addition to sampling highly relevant events to assist the agent in making accurate decisions, we randomly sample unrelated events to influence the agent’s behavior. For LLMs, we include random content in the prompt, which will have a certain impact on the output response. For the agent, we want it to behave like a real person, where spontaneous thoughts may influence its decision-making, or leading to a shift in the conversation topic. In our implementation, to prevent the frequent sampling of events that match the agent’s inner profile, we conduct weighted sampling based on the repetition rate. We employ the DBSCAN algorithm to cluster the embedding features of agent’s memory events $X=\{x_{1},x_{2},...,x_{n}\}$ into $k$ clusters denoted as $\{C_{1},C_{2},...,C_{k}\}$. For each cluster $C_{i}$, where $i=1,2,...,k$, we compute the number of events, $|C_{i}|$, contained within. The sampling probability $p_{i}$ for events in cluster $C_{i}$ is defined as: